A rectangular game board measures 126 centimeters by 90 centimeters. The game board is divided into small squares of equal size.
A) find the longest possible length of the side of one of the small square.
B) find the least number of squares the game board can be equally divided into.

Answer by ikleyn(53306)

.
A rectangular game board measures 126 centimeters by 90 centimeters. The game board is divided into small squares of equal size.

    A)  find the longest possible length of the side of one of the small square.      18 centimeters.      ANSWER

    B)  find the least number of squares the game board can be equally divided into.  7*5 = 35 squares.    ANSWER

Rationale :   because  126 = 18*7  and  90 = 18*5,  and  7  and  5  are relatively prime numbers.